The number of values of the pair (a,b) for which a(x+1)^2 + b{( – x^2) – 3x – 2} + x + 1 = 0 is an identity in x, is
1)0
2)1
3)2
4)Infinite
 

Grade:

1 Answers

Aditya Gupta
5 years ago
the eqn is:
(x+1)[ax+a-b(x+2)+1]=0
or x(a-b)+a-2b+1=0
or a=b and a-2b+1=0
so a=b=1
option 2) is correct